OpenStudy (anonymous):

if \(f(x)=ln~x\) and \(g(x)=9-x^2\), what is the domain of the \(f(g(x))\)?

OpenStudy (anonymous):

the domain of \(\ln(x)\) is \(x>0\) so your job is to solve \[9-x^2>0\]

OpenStudy (anonymous):

\[x^2<9~~?\]

OpenStudy (anonymous):

or wait..

OpenStudy (anonymous):

the domain of \(f(x)=ln~x\) is \(x>0\) the domain of \(g(x)=9-x^2\) is \(-3<x<3\) so that would make the domain of \(f(g(x))\) \(0<x<3\) yes? :D

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@Hero @ash2326

OpenStudy (ash2326):

@yummydum Why do you say domain of g(x) is bounded?

OpenStudy (anonymous):

aand i'm confused all over again...

hero (hero):

@yummydum, ask yourself, "Would something like -18 work for the domain?".

hero (hero):

If so, then you need to expand your domain.

OpenStudy (anonymous):

okay the domain for g(x) is \(x<\pm3\)?

OpenStudy (ash2326):

@yummydum There are no constraints for x in g(x), so the domain of g(x) should be?

OpenStudy (anonymous):

domain for f(x) 0<x domain for g(x) x<3 domain for f(g(x)) 0<x<3 ....I'm right, yea?

OpenStudy (anonymous):

there is a constraint because 9-x^2 has to be greater than 0 ...so x has to be less than 3

OpenStudy (ash2326):

Domain of g(x) is not correct, how do you arrive at this?

OpenStudy (ash2326):

Is that mentioned in the question, g(x) >= 0 ?

OpenStudy (anonymous):

o.O ....the equation says \[9-x^2\large\color{red}{>0}\]

OpenStudy (ash2326):

Nope, that's when you write f(g(x) \[f(g(x)=\ln (9-x^2)\] here, 9-x^2 >0 because logarithm can't have a negative no. or 0 inside it

OpenStudy (anonymous):

ok here we go \[9-x^2>0\] or \[(3-x)(3+x)>0\]

OpenStudy (anonymous):

the zeros are \(-3\) and \(3\) and since \[y=9-x^2\] is a parabola that opens down, it is positive between the zeros and negative outsides of them

OpenStudy (sirm3d):

@yummydum you found the answer, unfortunately you said it incorrectly with "the domain of g(x) is -3 < x < 3

OpenStudy (anonymous):

you want to know where it is positive, so you can take the log of the damn thing, so pick \[(-3,3)\]

OpenStudy (anonymous):

or if you prefer \[-3<x<3\]

OpenStudy (anonymous):

yea i know that part is wrong @sirm3d but the answer i said \(0<x<3\) is correct...yes?

OpenStudy (anonymous):

no it is not correct

OpenStudy (anonymous):

x-x

OpenStudy (sirm3d):

as @satellite73 said, "your job is to solve \(9-x^2>0\) "

OpenStudy (anonymous):

\(x\) for example could be \(-2\) because \(9-(-2)^2=9-4=5\)

OpenStudy (anonymous):

yes so x is less than 3

OpenStudy (anonymous):

yes?

OpenStudy (ash2326):

@yummydum we have \[9-x^2<0\] \[(3-x)(3+x)<0\] Let's draw number line

OpenStudy (anonymous):

okay..

OpenStudy (ash2326):

Let's choose a no. greater than 3, say 4 9-x^2 x=4 9-16=-7 obviously less than 0 So we know that x> 3 is not part of the solution Let's choose a no. less than -3 say =-4 9-x^2 x=-4 9-16=-7 again this also not a part of the solution. SO we have to check the region from -3 to 3. Do you follow this?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then what?

OpenStudy (ash2326):

ok, check x =-2 and x=1 Check if they work? 9-x^2 >0

OpenStudy (anonymous):

yes they do

OpenStudy (ash2326):

So that's your region or domain

OpenStudy (anonymous):

my choices are \[x\le3\]\[|x|\le3\]\[|x|>3\]\[|x|<3\]\[0<x<3\]

OpenStudy (ash2326):

ok if you have \[|x|>3\] \[=> x> 3\ or -x >3\] so \[x>3 \ or\ x<-3\] Is this match our answer?

OpenStudy (anonymous):

yes

OpenStudy (ash2326):

oh really? try putting x>3 in 9-x^2. Is that greater then 0 ?

OpenStudy (ash2326):

*than

OpenStudy (anonymous):

oh wait no...sorry got the signs mixed up

OpenStudy (ash2326):

then choose the correct answer, and you should have a reason for doing so

OpenStudy (anonymous):

\[|x|<3\] meaning x<3 and -x<3 -3<x<3 and thats the domain of 9-x^2

OpenStudy (ash2326):

yes, correct. Did you understand ?

OpenStudy (anonymous):

yes but the ln part confuses me..i thought it meant that x cannot be 0

OpenStudy (anonymous):

or wait...nvm in this case it means 9-x^2 cannot be 0

OpenStudy (ash2326):

yes, that's why only chose >0

hero (hero):

@ash, what happens if x = -18?

OpenStudy (anonymous):

i get it now, thanks @ash2326

OpenStudy (ash2326):

9-x^2=> -315 obviously negative no. can't be inside log

hero (hero):

Yeah, I forgot you have to square afterwards, sorry.

hero (hero):

@yummydum, that's not nice. I was only trying to help.

OpenStudy (ash2326):

@yummydum He was just trying to help, be polite and nice to other users.

OpenStudy (anonymous):

I mean stop as in thats confusing again..it took a while to register all that, wasn't trying to be rude

hero (hero):

I know I goofed up. Believe me, I'm cringing badly.

OpenStudy (ash2326):

ha ha Anyways @yummydum It's clear now, isn't it?

OpenStudy (anonymous):

yea, thanks

OpenStudy (anonymous):

another question really quick: \[\log_b\left({m\sqrt{n}\over p}\right)=\log_b m+{1\over2}\log_b n-\log_b p\]yes?

OpenStudy (ash2326):

yes, right

OpenStudy (anonymous):

okay and with this question am i doing it right?\[2\sin^2\theta=\sin\theta\]\[\sin\theta(2\sin\theta-1)=0\] \[sin\theta=0~~~~~~\sin\theta={1\over2}\]

OpenStudy (ash2326):

yes

OpenStudy (ash2326):

Anyways if you have other question, please post as a new one.

OpenStudy (anonymous):

thanks, will do